# Quasilinear elliptic systems: p-laplacian systems

### v1

@@ 4. The Concrete Palais-Smale condition
$\Omega_{1}^{\rho}=\{x\in\Omega\ :\ |u_{n}|\leq\rho\},$ $\Omega_{2}^{\rho}=\{x\in\Omega\ :\ |u_{n}|>\rho\}.$
@@
$\theta_{1,k}(s)=\begin{cases}s_{k}&|s|<\sigma\\ -s_{k}+2\sigma&\sigma\leq|s|\leq 2\sigma,\ s_{k}\geq 0\\ -s_{k}-2\sigma&\sigma\leq|s|\leq 2\sigma,\ s_{k}\leq 0\\ 0&|s|\geq\sigma.\end{cases}$
@@
Since $|\theta_{1,k}|\leq\sigma$, we get
@@
$=\frac{1}{\nu}\int_{\{|u_{n}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\frac{1}{\nu}\int_{\{\sigma\leq|u_{n}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
@@
$\int_{\{|u_{n}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-\int_{\{\sigma\leq|u_{n}|<2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad+\frac{1}{p_{k}}\int_{\{|u_{n}|\leq\sigma\}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}+\frac{1}{p_{k}}\int_{\{\sigma\leq|u_{n}|\leq 2\sigma\}}D_{s}A(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
@@
$\leq\tilde{C}(\sigma)+\frac{1}{\beta^{p_{k}^{\prime}}p_{k}^{\prime}}\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}^{p_{k}^{\prime}}+\frac{\beta^{p_{k}}}{\nu p_{k}}\int_{\{|u_{n}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad+\frac{\beta^{p_{k}}}{\nu p_{k}}\int_{\{\sigma\leq|u_{n}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
@@
$\frac{1}{p_{k}}\int_{\{|u_{n}|\leq\sigma\}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$ $+\frac{1}{p_{k}}\int_{\{\sigma\leq|u_{n}|\leq 2\sigma\}}D_{s}A(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
$\leq M\sigma\int_{\{|u_{n}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $+M\sigma\int_{\{\sigma\leq|u_{n}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Therefore, setting $I_{r}=\{|u_{n}|\leq r\}$ and $A_{r}=\{r\leq|u_{n}|\leq r+\sigma\}$:
@@
$\theta_{2,k}(s)=\begin{cases}0&|s|\leq\sigma\\ s_{k}-\sigma&\sigma<|s|<2\sigma,\quad s_{k}\geq 0\\ s_{k}+\sigma&\sigma<|s|<2\sigma,\quad s_{k}\leq 0\\ -s_{k}+3\sigma&2\sigma\leq|s|<3\sigma,\quad s_{k}\geq 0\\ -s_{k}-3\sigma&2\sigma\leq|s|<3\sigma,\quad s_{k}\leq 0\\ 0&|s|\geq 3\sigma,\end{cases}$
@@
$\theta_{k}(s)=\begin{cases}0&\text{$|s|\leq l\sigma$}\\ s_{k}-l\sigma&\text{$l\sigma<|s|<(l+1)\sigma$, $s_{k}\geq 0$}\\ s_{k}+l\sigma&\text{$l\sigma<|s|<(l+1)\sigma$, $s_{k}\leq 0$}\\ -\delta s_{k}-[1+\delta(l+1)]\sigma&\text{ $(l+1)\sigma\leq|s|<(l+1)\sigma+\frac{\sigma}{\delta}$, $s_{k}\leq 0$}\\ -\delta s_{k}+[1+\delta(l+1)]\sigma&\text{ $(l+1)\sigma\leq|s|<(l+1)\sigma+\frac{\sigma}{\delta}$, $s_{k}\geq 0$}\\ 0&\text{$|s|\geq(l+1)\sigma+\frac{\sigma}{\delta}$}.\end{cases}$

### v3

# On the multiplicity of weak solutions for a class of coupled quasilinear elliptic systems

@@ 4. The Concrete Palais-Smale condition
$\Omega_{1}^{\rho}=\{x\in\Omega\ :\ |u_{n,k}|\leq\rho\},$ $\Omega_{2}^{\rho}=\{x\in\Omega\ :\ |u_{n,k}|>\rho\}.$
@@
$\theta_{1,k}(s)=\begin{cases}s_{k}&|s_{k}|<\sigma\\ -s_{k}+2\sigma&\sigma\leq s_{k}\leq 2\sigma\\ -s_{k}-2\sigma&-2\sigma\leq s_{k}\leq-\sigma\\ 0&|s_{k}|\geq 2\sigma.\end{cases}$
@@
Since $|\theta_{1,k}|\leq\sigma$ and $\{u_{n,j}\}$ is uniformly bounded in $L^{q_{j}\frac{q_{k}-1}{q_{k}}}(\Omega)$, we get
@@
$=\frac{1}{\nu}\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\frac{1}{\nu}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
@@
$\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-\int_{\{\sigma\leq|u_{n,k}|<2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad+\frac{1}{p_{k}}\int_{\{|u_{n,k}|\leq\sigma\}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}+\frac{1}{p_{k}}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}D_{s}A(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
@@
$\leq\tilde{C}(\sigma)+\frac{1}{\beta^{p_{k}^{\prime}}p_{k}^{\prime}}\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}^{p_{k}^{\prime}}+\frac{\beta^{p_{k}}}{\nu p_{k}}\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad+\frac{\beta^{p_{k}}}{\nu p_{k}}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
@@
$\frac{1}{p_{k}}\int_{\{|u_{n,k}|\leq\sigma\}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$ $+\frac{1}{p_{k}}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}D_{s}A(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
$\leq M\sigma\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $+M\sigma\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Therefore, setting
$I_{r}=\{x\in\Omega:\ |u_{n,k}|\leq r\},\qquad A_{r}=\{x\in\Omega:\ r\leq|u_{n,k}|\leq r+\sigma\},$
we obtain
@@
$\theta_{2,k}(s)=\begin{cases}0&|s_{k}|\leq\sigma\\ s_{k}-\sigma&\sigma<s_{k}<2\sigma\\ s_{k}+\sigma&-2\sigma<s_{k}<-\sigma\\ -s_{k}+3\sigma&2\sigma\leq s_{k}<3\sigma\\ -s_{k}-3\sigma&-3\sigma\leq s_{k}<-2\sigma\\ 0&|s_{k}|\geq 3\sigma,\end{cases}$
@@
$\theta_{k}(s)=\begin{cases}0&\text{$|s_{k}|\leq l\sigma$}\\ s_{k}-l\sigma&\text{$l\sigma<s_{k}<(l+1)\sigma$}\\ s_{k}+l\sigma&\text{$-(l+1)\sigma<s_{k}<-l\sigma$}\\ -\delta\,s_{k}+\bigl[1+\delta(l+1)\bigr]\sigma,&(l+1)\sigma\leq s_{k}<(l+1)\sigma+\dfrac{\sigma}{\delta},\\ -\delta\,s_{k}-\bigl[1+\delta(l+1)\bigr]\sigma,&-(l+1)\sigma-\dfrac{\sigma}{\delta}<s_{k}\leq-(l+1)\sigma,\\ 0,&|s_{k}|\geq(l+1)\sigma+\dfrac{\sigma}{\delta}.\end{cases}$